座標の回転： R(θ) ( \begin{matrix}x \\ y \end{matrix} ) = ( \begin{matrix}cos θ & sin θ \\ − sin θ & cos θ \end{matrix} ) ( \begin{matrix}x \\ y \end{matrix} )

別の記述方法による出力例

座標の回転： R(θ) (\begin{matrix}x \\ y \end{matrix}) = (\begin{matrix}cos θ & sin θ \\ − sin θ & cos θ \end{matrix}) (\begin{matrix}x \\ y \end{matrix})